OpenStudy (anonymous):

Harold brought two leaves of raisin bread to share with his friends. If there are 12 people total, including Harold, how much of the bread does each person receive?

OpenStudy (anonymous):

2 loaves/12 people = 1/6 loaf each...definitely not enough raisin bread for anyone.

OpenStudy (anonymous):

These are the answer choices 0.06 0.16 0.25 0.3

OpenStudy (anonymous):

12 people 2 loaves divided into 12 portions each person gets 1/12th of each loaf x each person gets 1/12th of each loaf y each person gets 2/24th of total bread 1/12th of the total bread each

OpenStudy (anonymous):

0.16 one divided by 6

OpenStudy (anonymous):

yes, 6 people per loaf

OpenStudy (anonymous):

question is not worded that way though

OpenStudy (anonymous):

Thanks for your help! this question is confusing

OpenStudy (anonymous):

2 loaves divided into 12 portions--yes it is

OpenStudy (anonymous):

ya they each got 1/12 of each loaf, 1/12 + 1/12 = 1/6 darn im tired.

OpenStudy (anonymous):

12/6 = 2 loafs